# ADEdgeDrop: Adversarial Edge Dropping for Robust Graph Neural Networks

### v1

@@ V-A2 Compared Baselines and Backbones
In this paper, we compared the proposed ADEdgeDrop with representative state-of-the-art GAL methods, including edge-dropping, node-dropping, and adversarial training approaches. Herein, we present the code links and detailed descriptions of these methods as follows:
@@ V-A3 Implemented Details and Parameter Settings
The proposed ADEdgeDrop is implemented on a PyTorch platform with the Torch Geometric package. We run all experiments in this paper on a computer with AMD R9-5900HX, RTX 3080 16G GPU, and 32G RAM.
@@
There are four primary hyperparameters in our model, i.e., the threshold $\mu$, trade-off parameter $\alpha$, PGD iteration number $\eta$, and PGD updating rate $\gamma$. In all cases, the threshold $\mu$ ranges in $\{0.5,0.6,\cdots,1\}$. The PGD updating rate is fixed as $\eta=5$, and the other hyperparameters vary on different datasets. In our experiments, the values of both $\alpha$ and $\gamma$ are selected from $\{0.1,0.2,\cdots,1\}$. Detailed settings of these hyperparameters can be found in our codes.
@@
TABLE II: Performance comparison of various GAL methods for semi-supervised node classification, where the best accuracies are in bold and the second-best ones are underlined. OOT: out-of-time error.
@@
GNN Backbones GAL Methods Citeseer Pubmed ACM Chameleon UAI Actor
GCN Original 0.707 0.782 0.884 0.477 0.466 0.232
DropEdge [13] 0.716 0.784 0.888 0.465 0.548 0.251
DropNode [25] 0.713 0.783 0.887 0.476 0.552 0.242
GAUG-O [23] 0.719 OOT 0.891 0.469 0.502 0.241
FLAG [12] 0.709 0.789 0.889 0.498 0.500 0.266
DropMessage [26] 0.715 0.788 0.887 0.478 0.556 0.235
ADEdgeDrop 0.721 0.794 0.893 0.523 0.561 0.291
GAT Original 0.703 0.773 0.874 0.459 0.476 0.234
DropEdge [13] 0.708 0.775 0.879 0.467 0.539 0.237
DropNode [25] 0.705 0.772 0.881 0.469 0.535 0.261
GAUG-O [23] 0.706 OOT 0.892 0.483 0.496 0.259
FLAG [12] 0.706 0.778 0.876 0.489 0.502 0.236
DropMessage [26] 0.714 0.776 0.883 0.462 0.544 0.251
ADEdgeDrop 0.718 0.786 0.884 0.497 0.551 0.294
GraphSAGE Original 0.691 0.768 0.884 0.463 0.384 0.243
DropEdge [13] 0.692 0.770 0.890 0.472 0.474 0.268
DropNode [25] 0.707 0.778 0.887 0.483 0.431 0.265
GAUG-O [23] 0.728 OOT 0.888 0.413 0.442 0.249
FLAG [12] 0.698 0.777 0.885 0.498 0.481 0.245
DropMessage [26] 0.710 0.779 0.887 0.501 0.489 0.251
ADEdgeDrop 0.711 0.783 0.895 0.507 0.484 0.277
@@ V-B Experimental Results
V-B1 Performance with Different Backbones (RQ1).
@@
First, we evaluate the proposed ADEdgeDrop compared with representative baselines under different GNN backbones. For all GNNs, we use a 2-layer network architecture and fix the hidden dimension as 16. The learning rates are all fixed as 0.01 and the Adam optimizer is adopted. As reported in Table II, ADEdgeDrop obtains superior classification accuracy than other compared methods in most cases. For example, ADEdgeDrop achieves 8.47% and 2.54% improvements over the 2nd best methods on Actor and Chameleon datasets, respectively. GAUG-O which applies an edge predictor also displays competitive performance compared with other GAL methods. Nevertheless, it is time-consuming when addressing dense graphs with more edges, and even can not obtain the results in an acceptable time on Pubmed dataset. These observations indicate that ADEdgeDrop has a positive impact on robust representation learning compared with other state-of-the-art GAL methods, especially for random dropping methods.
Figure 3: GCN classification accuracy of different GAL methods with varying rates of addition edge attacks.
Figure 4: GCN classification accuracy of different GAL methods with varying rates of deletion edge attacks.
V-B2 Edge Attack (RQ2).
@@
In pursuit of testing the robustness of GNNs with the proposed method, we further conduct experiments on graphs with randomly added and removed edges. In detail, we randomly remove or add $5\%$ to $40\%$ edges of the original graphs as attacks while ensuring that the new corrupted graphs are still undirected graphs. We run each method 5 times and record the average performance of all compared methods. Note that GAUG-O is not able to cope with the Pubmed dataset in an acceptable time owing to the high time complexity. Figures 3 and 4 show the GCN classification performance with different GAL strategies when encountering distinct types of edge attacks with varied extents. The experimental results indicate that ADEdgeDrop succeeds in alleviating the performance decline caused by edge attacks. Generally, the edge addition is challenging to all GAL methods and it causes significant performance decline. This may owe to the disrupted correspondence information between nodes by the newly added edges, while deleted edges have less influence since they can often be replaced by other multi-hop connections. In most cases, ADEdgeDrop achieves competitive accuracy when the graphs are corrupted, demonstrating the effectiveness and robustness of our method when encountering edge attacks.
TABLE III: Percentage of deleted edges and GCN classification accuracy with incomplete graphs generated by the Random Dropping (RD) and ADEdgeDrop strategies.
Datasets Edge Deletion % RD ADEdgeDrop
Citeseer 32.03% 0.677 0.690
Pubmed 1.44% 0.785 0.787
ACM 14.97% 0.878 0.885
Chameleon 24.95% 0.471 0.485
V-B3 Training with Retained Edges (RQ2).
@@
Because the proposed approach tends to preserve critical connections during training, we further evaluate GCN with randomly removed edges and the retained edges that are stored in the optimal training iteration of ADEdgeDrop. Specifically, we train a new basic GCN with an incomplete graph learned by ADEdgeDrop. We also generate another incomplete graph by randomly removing the same number of edges. As reported in Table III, the percentage of deleted edges varies on distinct datasets because the optimal numbers of removed edges learned by ADEdgeDrop differ. It is noted that the experimental results reported in Table III sometimes are not as high as those in Table II. This is because we only use an incomplete graph for training in this experiment while results reported in Table II utilize a complete graph for testing and only remove some edges when training and optimizing parameters. Compared with results in Table II, GCN with the incomplete graph generated by ADEdgeDrop sometimes even outperforms GCN with the complete graph, e.g., on ACM and Chameleon datasets. This indicates that the learned graph may have higher quality than the original graph by removing some noisy edges. The quantitative results reveal that the model training with graphs learned by ADEdgeDrop gains superior accuracy than that with randomly deleted edges, which validates the feasibility of the incomplete graph learned by the proposed method.
V-B4 Ablation Study and Effectiveness (RQ3).
We next perform the ablation study to validate the effectiveness of the node similarity supervision, i.e., the effectiveness of the adversarial optimization target defined in Eq. (12). In light of this objective, we construct a simple edge predictor, which directly minimizes the training target $\mathcal{L}_{lg}$ defined in Eq. (15). Note that DropEdge in Table II can be regarded as a baseline without the edge predictor. Table IV reports the results of the ablation study, which reveals that ADEdgeDrop gains higher performance compared with that without the adversarial training. The experimental observation points out that the node attribute similarities are important to the proposed strategy, and the adversarial optimization objective has the ability to utilize node affinity information to explore more reasonable edge-dropping schema.
V-B5 Impact of Threshold $\mu$.
@@
We next conduct the experiments to investigate the influence of the threshold $\mu$, which determines the number of removed edges during training. We have the following observation from experimental results as demonstrated in Figure 5. First of all, we can find that more edges are deleted as the threshold $\mu$ increases. Note that $\mu=1$ does not necessarily mean that all edges have been removed due to the perturbation added to the output of the edge predictor. Second, on most datasets, the optimal accuracy is achieved when $\mu>0.5$, indicating that a suitable deletion rate of edges is beneficial to the node representation learning. On most datasets, it is significant that the accuracy tends to rise with the decreased number of edges until GNN reaches the highest accuracy. After that, the classification accuracy begins to fluctuate and dwindle, due to the excessive removal of edges in the graph. This indicates that the remaining neighbor information is insufficient to achieve satisfactory performance. On the UAI dataset, we obtain the best classification result when $\mu=1$, revealing that there exist a large number of redundant or noisy edges in this dataset. In conclusion, it is critical to set a suitable threshold $\mu$ in different datasets, similar to the dropout rate of random dropping methods.
@@
V-B6 Convergence Analysis.

### v2

@@ V-A2 Compared Baselines and Backbones
In this paper, we compare the proposed ADEdgeDrop with the state-of-the-art GAL methods, including edge-dropping, node-dropping, and adversarial training approaches. Herein, we present the code links and detailed descriptions of these methods as follows:
@@
  • DropEdge++ [^12] https://github.com/hanjq17/DropEdgePlus [35] integrates two structure-aware samplers, i.e., layer-increasingly-dependent sampler and the feature-dependent sampler, to drop edges during GNN training.
@@ V-A3 Implemented Details and Parameter Settings
The proposed ADEdgeDrop is implemented on a PyTorch platform with the Torch Geometric package. We run all experiments in this paper on a computer with AMD R9-5900X, RTX 4060Ti 16G GPU, and 32G RAM. Each algorithm is run 5 times and we record the average accuracy and standard deviation.
@@
There are four primary hyperparameters in our model, i.e., the threshold $\mu$, trade-off parameter $\alpha$, PGD iteration number $\eta$, and PGD updating rate $\gamma$. In all cases, the threshold $\mu$ ranges in $\{0.5,0.6,\cdots,1\}$. The PGD updating rate is fixed as $\eta=5$, and the other hyperparameters vary on different datasets. In our experiments, the values of both $\alpha$ and $\gamma$ are selected from $\{0.1,0.2,\cdots,1\}$. Note that we keep the consistency between backbone GNNs for the original graph and the line graph.
@@
TABLE II: Performance (Accuracyand Stdandard deviation%) comparison of various GAL methods for semi-supervised node classification, where the best accuracies are highlighted in bold and the second-best ones are underlined. OOT: Out-Of-Time (1 day) error, OOM: Out-Of-Memory (16G) error.
@@
GNN Backbones GAL Methods Metrics BlogCatalog Pubmed ACM Chameleon UAI Actor CoraFull Arxiv
GCN [5] Original Accuracy 70.64 77.14 87.84 48.21 56.92 20.76 56.94 69.61
Std $\pm$1.44 $\pm$0.44 $\pm$0.19 $\pm$0.26 $\pm$0.67 $\pm$1.88 $\pm$0.77 $\pm$0.15
DropEdge [17] Accuracy 71.22 77.24 88.94 48.34 57.82 19.94 57.99 69.83
Std $\pm$2.12 $\pm$0.93 $\pm$0.51 $\pm$0.94 $\pm$0.81 $\pm$0.48 $\pm$0.24 $\pm$0.48
DropNode [36] Accuracy 72.35 76.92 85.90 48.40 58.22 22.18 57.89 69.77
Std $\pm$1.2 $\pm$0.50 $\pm$0.41 $\pm$0.66 $\pm$1.74 $\pm$0.91 $\pm$0.35 $\pm$0.29
GAUG-O [33] Accuracy 70.89 OOT 88.80 49.12 57.03 22.90 OOM OOM
Std $\pm$0.67 $\pm$1.09 $\pm$1.11 $\pm$3.32 $\pm$0.44
FLAG [16] Accuracy 70.78 75.08 88.28 48.10 56.74 21.42 58.04 69.82
Std $\pm$0.75 $\pm$0.68 $\pm$0.35 $\pm$0.83 $\pm$1.66 $\pm$0.52 $\pm$0.33 $\pm$0.40
DropEdge++ [35] Accuracy 70.66 78.14 88.02 48.94 58.89 23.84 57.63 69.89
Std $\pm$3.32 $\pm$0.74 $\pm$0.39 $\pm$1.00 $\pm$0.98 $\pm$1.22 $\pm$1.14 $\pm$0.33
DropMessage [37] Accuracy OOM 78.49 87.62 49.22 58.70 21.94 57.04 OOM
Std $\pm$0.32 $\pm$2.13 $\pm$0.94 $\pm$1.98 $\pm$0.95 $\pm$2.35
ADEdgeDrop Accuracy 73.00 78.52 89.26 50.06 59.78 23.04 58.70 70.19
Std $\pm$0.68 $\pm$0.16 $\pm$0.08 $\pm$0.89 $\pm$0.81 $\pm$2.52 $\pm$0.46 $\pm$0.05
GAT [7] Original Accuracy 51.14 75.94 85.98 46.22 56.68 22.94 54.98 69.28
Std $\pm$5.34 $\pm$1.18 $\pm$0.50 $\pm$1.55 $\pm$1.30 $\pm$1.63 $\pm$0.59 $\pm$0.06
DropEdge [17] Accuracy 51.89 76.48 86.48 46.46 56.32 21.96 55.03 69.35
Std $\pm$4.13 $\pm$0.25 $\pm$0.42 $\pm$1.54 $\pm$0.93 $\pm$1.12 $\pm$0.47 $\pm$0.48
DropNode [36] Accuracy 52.91 76.18 85.92 46.56 48.78 22.40 55.79 69.53
Std $\pm$4.89 $\pm$0.35 $\pm$0.54 $\pm$1.76 $\pm$7.84 $\pm$1.87 $\pm$0.47 $\pm$0.39
GAUG-O [33] Accuracy 51.77 OOT 88.28 47.64 56.72 23.05 OOM OOM
Std $\pm$2.39 $\pm$0.45 $\pm$2.26 $\pm$5.10 $\pm$1.34
FLAG [16] Accuracy 48.44 76.30 87.16 48.08 58.04 21.80 57.38 69.76
Std $\pm$1.34 $\pm$0.46 $\pm$0.66 $\pm$1.27 $\pm$0.57 $\pm$0.72 $\pm$0.54 $\pm$0.31
DropEdge++ [35] Accuracy 53.22 76.72 87.21 47.56 57.87 23.69 57.34 69.63
Std $\pm$1.64 $\pm$0.76 $\pm$0.98 $\pm$0.73 $\pm$1.43 $\pm$1.23 $\pm$0.55 $\pm$0.22
DropMessage [37] Accuracy OOM 77.14 86.42 48.44 58.20 23.78 58.04 OOM
Std $\pm$0.46 $\pm$1.15 $\pm$0.90 $\pm$1.06 $\pm$1.54 $\pm$0.62
ADEdgeDrop Accuracy 54.52 76.86 87.86 48.60 58.58 25.60 57.70 69.70
Std $\pm$2.80 $\pm$1.03 $\pm$0.52 $\pm$1.19 $\pm$0.80 $\pm$1.37 $\pm$0.41 $\pm$0.10
GraphSAGE [6] Original Accuracy 75.30 77.30 86.16 47.12 55.20 23.04 57.34 69.28
Std $\pm$2.87 $\pm$0.81 $\pm$0.47 $\pm$1.70 $\pm$0.95 $\pm$1.63 $\pm$0.67 $\pm$0.11
DropEdge [17] Accuracy 76.43 76.98 87.52 48.54 57.48 23.22 58.79 69.44
Std $\pm$3.47 $\pm$1.95 $\pm$0.25 $\pm$0.99 $\pm$0.84 $\pm$0.62 $\pm$0.72 $\pm$0.20
DropNode [36] Accuracy 77.02 76.28 87.62 47.96 56.94 21.86 57.98 69.73
Std $\pm$2.82 $\pm$1.76 $\pm$0.43 $\pm$1.04 $\pm$1.92 $\pm$0.45 $\pm$0.44 $\pm$0.29
GAUG-O [33] Accuracy 77.45 OOT 88.10 47.36 55.77 23.76 OOM OOM
Std $\pm$3.50 $\pm$0.33 $\pm$0.83 $\pm$6.04 $\pm$1.04
FLAG [16] Accuracy 75.70 74.04 86.14 48.12 57.48 22.22 57.84 69.62
Std $\pm$0.88 $\pm$1.29 $\pm$0.76 $\pm$1.08 $\pm$1.67 $\pm$0.98 $\pm$0.50 $\pm$0.16
DropEdge++ [35] Accuracy 78.61 77.52 87.93 48.31 58.11 23.49 58.71 69.58
Std $\pm$1.54 $\pm$0.44 $\pm$0.67 $\pm$0.65 $\pm$1.98 $\pm$2.11 $\pm$0.85 $\pm$0.38
DropMessage [37] Accuracy OOM 78.08 86.48 47.12 59.40 21.76 57.58 OOM
Std $\pm$0.39 $\pm$0.15 $\pm$0.53 $\pm$1.74 $\pm$2.37 $\pm$2.56
ADEdgeDrop Accuracy 80.30 78.12 88.36 48.68 62.98 25.06 58.98 70.35
Std $\pm$2.05 $\pm$0.65 $\pm$0.41 $\pm$0.93 $\pm$1.21 $\pm$2.25 $\pm$0.62 $\pm$0.12
@@ V-B Experimental Results
V-B1 Performance with Different Backbones (RQ1)
@@
First, we evaluate the proposed ADEdgeDrop compared with representative baselines under different GNN backbones. The learning rates of all models are fixed as 0.01 and the Adam optimizer is adopted. For all GNN backbones except those with DropEdge++, we use a 2-layer network architecture and fix the hidden dimension as 16 on all datasets except Arxiv. We fix the hidden dimension of backbones on Arxiv dataset as 256 for all models. Moreover, we also adopt the pre-dropping process on the edge-dense Arxiv dataset. As reported in Table II, ADEdgeDrop obtains superior classification accuracy than other compared methods in most cases. Notably, to achieve desired performance, we need to fix the dimension of hidden layers as 128 or even more for DropEdge++ on BlogCatalog and UAI datasets, which are much larger than all other methods. This indicates DropEdge++ requires more computational resources to achieve comparable accuracy. In addition, GAUG-O which also adopts an edge predictor exhibits competitive performance compared with other GAL methods. Nevertheless, it is time-consuming when addressing dense graphs with more edges, and even cannot obtain the results in an acceptable time on Pubmed, CoraFull and Arxiv datasets. Compared with FLAG, which conducts adversarial training on features, the proposed ADEdgeDrop generally shows promoting performance. This observation indicates that adversarial edge learning improves GNN training more significantly. In conclusion, the experimental results indicate that ADEdgeDrop positively impacts robust representation learning compared with other state-of-the-art GAL methods, especially for random dropping methods like DropEdge, DropNode, DropEdge++ and DropMessage.
V-B2 Performance with Edge-Removal and Edge-Addition Attacks (RQ2)
@@
In pursuit of testing the robustness of GNNs with the proposed method, we further conduct experiments on graphs with randomly added and removed edges. In detail, we randomly remove or add $5\%$ to $40\%$ edges of the original graphs as attacks while ensuring that the new corrupted graphs are still undirected graphs. Note that GAUG-O is not able to cope with the Pubmed dataset in an acceptable time owing to the high time complexity. Figures 3 and 4 show the GCN classification performance with different GAL strategies when encountering distinct types of edge-removal or edge-addition attacks with varied extents. The experimental results indicate that ADEdgeDrop succeeds in alleviating the performance decline caused by edge-removal or edge-addition attacks. Generally, the edge addition is more challenging to all GAL methods and it causes significant performance decline. This may owe to the disrupted correspondence information between nodes by the newly added edges, while deleted edges have less influence since other multi-hop connections can often replace them. In most cases, ADEdgeDrop achieves competitive accuracy when the graphs are corrupted, demonstrating the effectiveness and robustness of our method when encountering edge-removal or edge-addition attacks.
Figure 3: GCN classification accuracy of different GAL methods with varying rates of edge-addition attacks.
Figure 4: GCN classification accuracy of different GAL methods with varying rates of edge-removal attacks.
V-B3 Analysis on Edge Dropping (RQ2)
First, we generate a toy example to visualize the edge dropping results intuitively, as shown in Figure 5. There are 971 edges and 2 classes of nodes in this toy graph, where the pink edges indicate the noisy connections in the graph. We visualize the graphs learned by the proposed ADEdegeDrop and other edge-based methods, from which we have the following observations. We set the widely used drop rate for the random edge-dropping method, i.e., DropEdge, as 0.5. It can be found that DropEdge gets a sparser graph with higher classification accuracy. GAUG-O resamples the edges, thereby adding more connections and removing some unnecessary relationships simultaneously, leading to better performance. As for the proposed ADEdgeDrop, our method significantly removes more noisy connections marked in pink. It can be verified that ADEdgeDrop improves the classification results and obtains a high-quality graph with less noisy or redundant connections. These experiments results also validate that most edges in a graph may be redundant, and the GNN training on an incomplete graph often leads to competitive performance.
Figure 5: Visualization of retained edges generated by DropEdge, GAUG-O and ADEdgeDrop on the toy example dataset, where pink edges indicate noisy connections.
@@ V-B3 Analysis on Edge Dropping (RQ2)
In light of previous experimental analysis, a large number of edges in a graph are unnecessary or alternative. With the message passing mechanism of GNNs, some edges can also be replaced by other multi-order paths. Thus, it is reasonable to learn a sparse and pure graph where important connections are retained and some redundant relationships are removed. Because the proposed model tends to preserve critical connections during training, we further evaluate GCN with randomly removed edges and the retained edges that are stored in the optimal training iteration of ADEdgeDrop. In detail, we train a new basic GCN with an incomplete graph learned by ADEdgeDrop. We also generate another incomplete graph by randomly removing the same number of edges. As reported in Table III, the percentage of deleted edges varies on distinct datasets because the optimal numbers of removed edges learned by ADEdgeDrop differ. It is noted that the experimental results on graph $\hat{\mathcal{G}}_{AD}$ reported in Table III are not as high as the ADEdgeDrop model in Table II. This is because we only use an incomplete graph for training in this experiment, while results reported in Table II utilize a complete graph for testing and only remove some edges when training and optimizing parameters. However, compared with the performance of GCN with complete graph $\mathcal{G}$, GCN with the incomplete graph $\hat{\mathcal{G}}_{AD}$ generated by ADEdgeDrop still obtains comparable performance. On BlogCatalog, ACM and Actor datasets, GCN with $\hat{\mathcal{G}}_{AD}$ even gains higher accuracy. This indicates that the learned graph may have higher quality than the original graph by removing some noisy edges, and in most cases the backbone GCN can still achieve competitive performance with sparser graphs learned by ADEdgeDrop. In addition, the quantitative results also reveal that the model training with graphs $\hat{\mathcal{G}}_{AD}$ learned by ADEdgeDrop gains superior accuracy than that training with $\hat{\mathcal{G}}_{RD}$ generated by random edge-dropping method, thereby validating the feasibility of the incomplete graph learned by the proposed method.
TABLE III: Percentage of deleted edges ($\varsigma$) and GCN classification accuracy (%) with complete graphs ($\mathcal{G}$), incomplete graphs generated by the random edge dropping ($\hat{\mathcal{G}}_{RD}$) and ADEdgeDrop strategies ($\hat{\mathcal{G}}_{AD}$). The best accuracies are highlighted in bold and the second-best ones are underlined.
Datasets $\varsigma$ GCN + $\mathcal{G}$ GCN + $\hat{\mathcal{G}}_{RD}$ GCN + $\hat{\mathcal{G}}_{AD}$
BlogCatalog 32.0 70.64 $\pm$ 1.44 70.13 $\pm$ 0.94 70.68 $\pm$ 0.65
Pubmed 56.9 77.14 $\pm$ 0.44 72.22 $\pm$ 0.80 74.81 $\pm$ 0.71
ACM 69.7 87.84 $\pm$ 0.19 85.14 $\pm$ 1.00 88.93 $\pm$ 0.10
Chameleon 49.6 48.21 $\pm$ 0.26 44.92 $\pm$ 0.85 47.23 $\pm$ 0.40
UAI 57.1 56.92 $\pm$ 0.67 54.33 $\pm$ 0.31 56.64 $\pm$ 0.54
Actor 22.0 20.76 $\pm$ 1.88 21.16 $\pm$ 1.39 23.71 $\pm$ 0.93
CoraFull 35.0 56.94 $\pm$ 0.77 54.62 $\pm$ 1.13 55.95 $\pm$ 0.99
Arxiv 24.8 69.61 $\pm$ 0.15 64.01 $\pm$ 0.17 68.22 $\pm$ 0.04
V-B4 Impact of Threshold $\mu$ (RQ2)
@@
We next conduct the experiments to investigate the influence of the threshold $\mu$, which determines the number of removed edges during training. We have the following observations from experimental results as demonstrated in Figure 6. First of all, we can find that more edges are deleted as the threshold $\mu$ increases. Note that $\mu=1$ does not necessarily mean that all edges have been removed due to the perturbation added to the output of the edge predictor. Second, on most datasets, the optimal accuracy is achieved when $\mu>0.5$, indicating that a suitable deletion rate of edges is beneficial to the node representation learning. On most datasets, it is significant that the accuracy tends to rise with the decreased number of edges until GNN reaches the highest accuracy. After that, the classification accuracy may begin to fluctuate and dwindle, due to the excessive removal of edges in the graph. This indicates that the remaining neighbor information is insufficient to achieve satisfactory performance. In conclusion, it is critical to set a suitable threshold $\mu$ in different datasets, similar to the dropout rate of random dropping methods.
@@
TABLE IV: Ablation study of ADEdgeDrop with GCN as the backbone. ADEdgeDrop w/o AT: ADEdgeDrop without the Adversarial Training. The best accuracies are highlighted in bold.
Datasets ADEdgeDrop w/o AT ADEdgeDrop
BlogCatalog 71.94 $\pm$ 0.68 73.00 $\pm$ 0.68
Pubmed 78.00 $\pm$ 0.71 78.52 $\pm$ 0.16
ACM 88.82 $\pm$ 0.15 89.26 $\pm$ 0.08
Chameleon 36.78 $\pm$ 3.50 50.06 $\pm$ 0.89
UAI 57.62 $\pm$ 1.05 59.78 $\pm$ 0.81
Actor 22.40 $\pm$ 1.45 23.04 $\pm$ 2.52
CoraFull 57.50 $\pm$ 0.65 58.70 $\pm$ 0.46
Arxiv 69.43 $\pm$ 0.13 70.19 $\pm$ 0.05
V-B5 Ablation Study and Effectiveness (RQ3)
We further perform the ablation study to validate the effectiveness of the node similarity supervision, i.e., the effectiveness of the adversarial optimization target defined in Eq. (12). In light of this objective, we construct a simple edge predictor, which directly minimizes the training target $\mathcal{L}_{lg}$ defined in Eq. (15). Note that DropEdge in Table II can be regarded as a baseline without the edge predictor. Table IV reports the results of the ablation study, which reveals that ADEdgeDrop gains higher performance compared with that without the adversarial training. The experimental observation points out that the node attribute similarities are important to the proposed strategy, and the adversarial optimization objective has the ability to utilize node affinity information to explore more reasonable edge-dropping schema.
V-B6 Training Convergence Analysis (RQ3)